# Locally rigid $\infty$-categories

### v1

@@ 4.2 Rigidity
In the locally rigid case, we have seen in 4.21 (although not proved yet) that the duality data of $\mathcal{W}$ can be described in terms of the multiplication map and an “exceptional” functor. It turns out that in the rigid case, this exceptional functor is simply the global sections functor, so that one can define the duality data “ahead of time”. In particular, one may deduce dualizability without assuming it. More precisely:
@@
[…]
@@
[…]
@@
Thus, by [Ram24, Addendum 2.42], the multiplication map is an internal left adjoint in $\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}}$. If we know that $S$-telescopes generate $\mathcal{W}$ under colimits, 3.7 then directly implies that the right adjoint to the multiplication is $\mathcal{W}$-linear on either side, and thus $\mathcal{W}$ is rigid .
@@
From all of these niceness properties, one finds that when working over a rigid base, most of the theory developed over $\mathbf{Sp}$ works essentially the same. In particular, the theory of localizing invariants (which can be defined over any base) behaves quite closely to the classical theory over rigid bases. For example, a consequence of 4.46 is that ordinary $K$-theory of the underlying (dualizable) $\mathcal{V}$-module is the universal localizing invariant on $\mathcal{V}$-modules. Note that this is not true over a general base. In fact, a question worth investigating could be:
Question 4.58.
Let $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$. What is the analogue of $K$-theory on $\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}}$
@@
As $\mathcal{W}$ is rigid over $\mathcal{V}$, $\iota$ is also a $\mathcal{W}$-linear left adjoint (we explain this in the lemma following the proof). Since $\mathcal{W}$ is rigid over $\mathcal{V}$, it suffices to prove that $\mathcal{W}_{L}$ is locally rigid over $\mathcal{W}$ by Example 4.12, and thus we are reduced without loss of generality to the case $\mathcal{V}=\mathcal{W}$.
@@ 4.3 Rigidification
Our goal in this section is to introduce rigidifications. Beyond the fact that this is a very natural notion, we are mainly after the following fact: every locally rigid $\mathcal{V}$-algebra is the “completion” of a rigid $\mathcal{V}$-algebra, namely its rigidification. This is convenient, as it allows us to reduce many properties of locally rigid categories to analogous ones for rigid categories - particularly, it will allow us to give an alternative proof of 4.21 and to get a more precise understanding of the exceptional functor $\Gamma_{!}$ appearing there.
@@
Recall that by 4.59, if this canonical map admits a fully faithful left adjoint, then $\mathcal{W}$ is locally rigid.
@@
We will prove this theorem by “describing” the rigidification in a similar fashion to the dualizable core from 2.44. Before doing so, let us explain the main consequence, that is, the promised alternative proof of 4.21:
@@
[…]
@@
By 4.60, $p^{L}$ is $\overline{\mathcal{W}}$-linear, so that the canonical map $p^{L}(w\otimes p(\overline{w}))\to p^{L}(w)\otimes\overline{w}$ is an equivalence for all $w\in\mathcal{W},\overline{w}\in\overline{\mathcal{W}}$.
@@
Altogether, this simply proves that the pre-duality datum we provided is exactly the one induced by the one of $\overline{\mathcal{W}}$ from 4.54 using the retraction $(p^{L},p)$, and hence is itself a duality datum. ∎
@@
In this case, $\mathcal{W}$ is locally rigid by 4.59 and the exceptional functor $\Gamma_{!}:\mathcal{W}\to\mathcal{V}$ is given by ${\mathrm{at}}_{\mathcal{W}}(\mathbf{1}_{\mathcal{W}},-)$.
@@
In particular, $p^{L}$ is a fully faithful internal left adjoint, so that by 2.21 we have an equivalence ${\mathrm{at}}_{\mathcal{W}}(x,y)\simeq{\mathrm{at}}_{\overline{\mathcal{W}}}(p^{L}(x),p^{L}(y))$. Since $\overline{\mathcal{W}}$ is rigid, the latter is
@@
Now, $p$ makes $\mathcal{W}$ into a $\overline{\mathcal{W}}$-module, and since $p^{L}$ is $\mathcal{V}$-linear and $\overline{\mathcal{W}}$ is rigid over $\mathcal{W}$, $p^{L}$ is $\overline{\mathcal{W}}$-linear by 4.60. It follows then that $\hom^{\overline{\mathcal{W}}}_{\overline{\mathcal{W}}}(p^{L}(x),\mathbf{1}_{\overline{\mathcal{W}}})\simeq\hom^{\overline{\mathcal{W}}}_{\mathcal{W}}(x,p(\mathbf{1}_{\overline{\mathcal{W}}}))=\hom^{\overline{\mathcal{W}}}_{\mathcal{W}}(x,\mathbf{1}_{\mathcal{W}})=p^{R}\hom_{\mathcal{W}}^{\mathcal{W}}(x,\mathbf{1}_{\mathcal{W}})$ and so with $x=\mathbf{1}_{\mathcal{W}}$, we find $\hom_{\overline{\mathcal{W}}}^{\overline{\mathcal{W}}}(p^{L}(\mathbf{1}_{W}),\mathbf{1}_{\overline{\mathcal{W}}})=p^{R}(\mathbf{1}_{\mathcal{W}})$.
@@
The final thing we do before describing rigidifications in order to prove 4.63 is to prove that they exist.
@@
For this, we simply note that since the unit of $\mathcal{W}$ is $\kappa$-compact, any symmetric monoidal functor $\mathcal{W}_{0}\to\mathcal{W}$ has a $\kappa$-filtered colimit-preserving right adjoint by 4.49. Thus we may replace the right hand side by $\operatorname{Fun}^{L,\otimes,\kappa}_{\mathcal{V}}(\mathcal{W}_{0},\mathcal{W})$.
Similarly, the unit of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ is atomic, and so by 4.48, we may replace the left hand side by $\operatorname{Fun}^{iL,\otimes}_{\mathcal{V}}(\mathcal{W}_{0},\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa}))$.
@@
By 4.67, it suffices to prove that $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ has a rigidification.
Consider the poset $P$ of full-sub-$\mathcal{V}$-algebras of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ that are rigid. By 4.48, for $\mathcal{W}_{0}\in P$, the inclusion $\mathcal{W}_{0}\to\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ is an internal left adjoint and in particular preserves $\kappa$-compacts, and $\mathcal{W}_{0}$ is $\kappa$-compactly generated - thus $P$ is small : it is a subposet of the poset of subcategories of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})^{\kappa}$.
@@
Let $\mathcal{W}_{0}:=\mathrm{colim}_{C\in P}C$, the colimit being taken in $\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}))$. By 4.53, $\mathcal{W}_{0}$ is rigid, and it certainly has a map to $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$. Its image in the sense of 4.68 is rigid (by that lemma) and contains $C$ for every $C\in P$, thus proving that it has a largest element, thereby concluding the proof. ∎
@@ 4.3.1 Over $\mathbf{Sp}$
In the same vein as for [Ram24, Section 4], while the existence of rigidification is good to know, for certain purposes it can be useful to have a more concrete description (for example, to prove 4.63). Our goal here is to draw inspiration from 4.74 to construct the rigidification by hand in terms of trace-class maps.
@@
This follows from 4.73: in a locally rigid category, every compact map is trace-class by 4.15. ∎
@@
(Rigidity) We first prove that $\mathrm{Rig}_{\kappa}(\mathcal{V})$ is rigid. Our goal is to apply 4.56. First, the unit is $y(\mathbf{1}_{\mathcal{V}})$ and hence compact.
@@
It follows from 4.56 that $\mathrm{Rig}_{\kappa}(\mathcal{V})$ is rigid.
@@
are all equivalences. The first and last one are equivalences because $\mathcal{M}$ is rigid and the unit of both $\mathrm{Rig}_{\kappa}(\mathcal{V}),\mathrm{Ind}(\mathcal{V}^{\kappa})$ are compact, so that 4.48 applies. For the middle one, we note that by 4.73 and 4.15, $\mathcal{M}$ is generated under colimits by $\mathbb{Q}_{\geq 0}$-indexed diagrams along trace-class maps, so it suffices to show that these land in $\mathrm{Rig}_{\kappa}(\mathcal{V})$ for any symmetric monoidal internal left adjoint $f:\mathcal{M}\to\mathrm{Ind}(\mathcal{V}^{\kappa})$.
@@
Now let $m\in\mathcal{M}$ be compactly exhaustible, so we write $m\simeq\mathrm{colim}_{\mathbb{N}}m_{n}$ with trace-class transition maps by 4.15. Each of the $f(m_{n})\to f(m_{n+1})$ factors through a $y(z_{n}),z_{n}\in\mathcal{V}^{\kappa}$ by the above observation, where $f(m_{n})\to y(z_{n})$ is trace-class. Thus $f(m)\simeq\mathrm{colim}_{\mathbb{N}}y(z_{n})$ with each $z_{n}\to z_{n+1}$ trace-class.
@@
where the inclusion $[1]\to[2]$ is the long edge, and where $\mathbb{Q}_{\geq 0}^{\delta}$ denotes the set $\mathbb{Q}_{\geq 0}$ viewed as a discrete category.
@@
[…]
@@
As $P$ is linearly ordered, it is a filtered colimit of posets of the form $[n]$, and so we may assume without loss of generality that it is of this form. Since the pushouts $[n]\coprod_{n=0}[1]\simeq~{}[n+1]$ are preserved by $(-)^{\simeq}$ (noting that $P^{\simeq}\simeq P^{\delta}$), we may also assume $P=[1]$.
@@
1. was proved in 4.71.
2. If $\mathcal{V}$ is locally rigid, then compact maps are trace-class, and it follows that $\hat{y}:\mathcal{V}\to\mathrm{Ind}(\mathcal{V}^{\kappa})$ lands in $\mathrm{Rig}(\mathcal{V})$ by 4.73 : indeed, compactly exhaustible objects generate $\mathcal{V}$ under colimits, and if $x=\mathrm{colim}_{\mathbb{Q}_{\geq 0}}x_{r}$ with compact transition maps, then $\hat{y}(x)=\mathrm{colim}_{\mathbb{Q}_{\geq 0}}y(x_{r})$, and each $y(x_{r})\to y(x_{s})$ is trace-class, as it was so in $\mathcal{V}$. In particular, $\hat{y}$ the desired fully faithful left adjoint.
Conversely, if there is such a left adjoint, then 4.59 kicks in, showing that $\mathcal{V}$ is locally rigid. ∎
@@
Construction 4.82.
Let $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$. By Example 4.35, $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})$ is rigid and hence the canonical symmetric monoidal colimit-preserving functor $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})\to\mathcal{V}$ factors through $\mathrm{Rig}(\mathcal{V})$. Beware that in general, the induced map $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})\to\mathrm{Rig}(\mathcal{V})$, while always fully faithful, is not in general an equivalence.
@@
Indeed consider, in $\prod_{i}\mathcal{V}_{i}$, a trace-class map $(v_{i})\to(w_{i})$. Each $v_{i}\to w_{i}$ is trace-class in $\mathcal{V}_{i}$, and hence compact ($\mathbf{1}_{\mathcal{V}_{i}}$ is compact) and hence factors through a compact ($\mathcal{V}_{i}$ is compactly generated) and hence through a dualizable ($\mathcal{V}_{i}$ is rigid). As dualizables in $\prod_{i}\mathcal{V}_{i}$ are detected pointwise, this implies that $(v_{i})\to(w_{i})$ factors through a dualizable object. Thus the result follows from the following example.
More generally, as an application of this description of rigidification together with 4.27, we obtain:
@@
Note that this forgetful functor exists by 4.48. Before proving this, we will need a preliminary which is about trace-class maps in lax pullbacks. First, we analyze internal homs in lax pullbacks, by analyzing a simpler case.
@@
Since $y\to z$ is also trace-class in $C_{01}$, 4.30 guarantees that there is a homotopy between the induced lifts for $gf$, and thus that $gf$ admits a trace-class witness in $C$, as was needed. ∎
@@
By 4.86, the span $\overset{\rightarrow}{\mathrm{fib}}(p_{0})\to C_{01}\leftarrow\overset{\rightarrow}{\mathrm{fib}}(p_{1})$ fits the hypotheses of the previous corollary. It thus suffices to show that the projection of $f$ (resp. $g$) to $\overset{\rightarrow}{\mathrm{fib}}(p_{i})$ is trace-class.
@@
${{f(x)}}$${{g(x)}}$${{f(y)}}$${{g(y)}}$$\scriptstyle{\eta_{x}}$$\scriptstyle{f(\alpha)}$$\scriptstyle{g(\alpha)}$$\scriptstyle{\eta_{y}}$
From there, using 4.74 and the fact that $f,g$ commute with sequential colimits, we find that $\eta$ is an equivalence on $\omega_{1}$-compacts and thus on all objects because $f,g$ commute with all colimits. ∎
@@
Let $\mathcal{V}_{0}\xrightarrow{p_{0}}\mathcal{V}_{01}\xleftarrow{p_{1}}\mathcal{V}_{1}$ be span in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$, and let $\mathcal{V}$ be their pullback, $\overset{\rightarrow}{\mathcal{V}}$ be their lax pullback as in 4.89. The canonical inclusion $\mathcal{V}\to\overset{\rightarrow}{\mathcal{V}}$ induces an equivalence on rigidification.
@@
We are finally equipped to give a proof of 4.85:
Proof of 4.85.
By definition of rigidity, this forgetful functor factors as a composite $\mathrm{CAlg}^{\mathrm{rig}}\to\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})\to(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}}$ where the second functor preserves limits for general reasons; thus we are really considering the first functor $\mathrm{CAlg}^{\mathrm{rig}}\to\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})$. Now, again by 4.48 this is a full subcategory inclusion, so it suffices to prove that it is closed under limits.

### v2

@@ 4.2 Rigidity
In the locally rigid case, we have seen in Section˜4.1 (although not proved yet) that the duality data of $\mathcal{W}$ can be described in terms of the multiplication map and an “exceptional” functor. It turns out that in the rigid case, this exceptional functor is simply the global sections functor, so that one can define the duality data “ahead of time”. In particular, one may deduce dualizability without assuming it. More precisely:
@@
See [HSS+21, Proposition 2.17]. The point is that the following functors provide a (self-)duality datum for $\mathcal{W}$: $\mathcal{V}\xrightarrow{\eta}\mathcal{W}\xrightarrow{\mu^{R}}\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}$ and $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\xrightarrow{\mu}\mathcal{W}\xrightarrow{\eta^{R}}\mathcal{V}$. Checking the triangle identities will simply use the $\mathcal{W}$-linearity of $\mu^{R}$ and the $\mathcal{V}$-linearity of $\eta^{R}$ - the details are in the cited reference. ∎
@@
Let $\mathcal{W}$ be a rigid commutative $\mathcal{V}$-algebra. The functor $\mathcal{W}\to\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{W},\mathcal{V})$ given as the mate of $\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\xrightarrow{\mu}\mathcal{W}\xrightarrow{\eta^{R}}\mathcal{V}$ is an equivalence. In particular, any $\mathcal{V}$-linear functor $\mathcal{W}\to\mathcal{V}$ is of the form $\hom_{\mathcal{W}}(\mathbf{1}_{\mathcal{W}},x\otimes-)$ for a uniquely determined $x$.
@@
Thus, by [RAM24, Addendum 2.42], the multiplication map is an internal left adjoint in $\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}}$. If we know that $S$-telescopes generate $\mathcal{W}$ under colimits, Section˜3 then directly implies that the right adjoint to the multiplication is $\mathcal{W}$-linear on either side, and thus $\mathcal{W}$ is rigid .
@@
From all of these niceness properties, one finds that when working over a rigid base, most of the theory developed over $\mathbf{Sp}$ works essentially the same. In particular, the theory of localizing invariants (which can be defined over any base) behaves quite closely to the classical theory over rigid bases. For example, a consequence of Section˜4.2 is that ordinary $K$-theory of the underlying (dualizable) $\mathcal{V}$-module is the universal localizing invariant on $\mathcal{V}$-modules. Note that this is not true over a general base. In fact, a question worth investigating could be:
Question 4.61.
Let $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$. What is the analogue of $K$-theory on $\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}}$ ?
@@
As $\mathcal{W}$ is rigid over $\mathcal{V}$, $\iota$ is also a $\mathcal{W}$-linear left adjoint (we explain this in the lemma following the proof). Since $\mathcal{W}$ is rigid over $\mathcal{V}$, it suffices to prove that $\mathcal{W}_{L}$ is locally rigid over $\mathcal{W}$ by Section˜4.1, and thus we are reduced without loss of generality to the case $\mathcal{V}=\mathcal{W}$.
@@ 4.3 Rigidification
Our goal in this section is to introduce rigidifications. Beyond the fact that this is a very natural notion, we are mainly after the following fact: every locally rigid $\mathcal{V}$-algebra is the “completion” of a rigid $\mathcal{V}$-algebra, namely its rigidification. This is convenient, as it allows us to reduce many properties of locally rigid categories to analogous ones for rigid categories - particularly, it will allow us to give an alternative proof of Section˜4.1 and to get a more precise understanding of the exceptional functor $\Gamma_{!}$ appearing there.
@@
The following is an obvious observation, but worth making nonetheless:
Observation 4.66.
Let $\mathcal{V}_{0}\to\mathcal{V}_{1}\to\mathcal{W}$ be maps in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$, and suppose $\mathcal{V}_{1}$ is $\mathcal{V}_{0}$-rigid. In this case, $\mathcal{W}$ admits a $\mathcal{V}_{0}$-rigidification if and only if it admits a $\mathcal{V}_{1}$-rigidification, and in case they exist, the canonical map between them is an equivalence.
Proof.
This follows by considering the respective universal properties, and using Section˜4.1 and Section˜4.1. ∎
@@
Recall that by Section˜4.2, if this canonical map admits a fully faithful left adjoint, then $\mathcal{W}$ is locally rigid.
@@
We will prove this theorem by “describing” the rigidification in a similar fashion to the dualizable core from Section˜2. Before doing so, let us explain the main consequence, that is, the promised alternative proof of Section˜4.1:
@@
$\mathcal{V}\xrightarrow{\eta}\mathcal{W}\xrightarrow{\mu^{R}}\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W},\qquad\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\xrightarrow{\mu}\mathcal{W}\xrightarrow{\Gamma_{!}}\mathcal{V}$
@@
By Section˜4.2, $p^{L}$ is $\overline{\mathcal{W}}$-linear, so that the canonical map $p^{L}(w\otimes p(\overline{w}))\to p^{L}(w)\otimes\overline{w}$ is an equivalence for all $w\in\mathcal{W},\overline{w}\in\overline{\mathcal{W}}$.
@@
Altogether, this simply proves that the pre-duality datum we provided is exactly the one induced by the one of $\overline{\mathcal{W}}$ from Section˜4.2 using the retraction $(p^{L},p)$, and hence is itself a duality datum. ∎
@@
In this case, $\mathcal{W}$ is locally rigid by Section˜4.2 and the exceptional functor $\Gamma_{!}:\mathcal{W}\to\mathcal{V}$ is given by ${\mathrm{at}}_{\mathcal{W}}(\mathbf{1}_{\mathcal{W}},-)$.
@@
In particular, $p^{L}$ is a fully faithful internal left adjoint, so that by Section˜2 we have an equivalence ${\mathrm{at}}_{\mathcal{W}}(x,y)\simeq{\mathrm{at}}_{\overline{\mathcal{W}}}(p^{L}(x),p^{L}(y))$. Since $\overline{\mathcal{W}}$ is rigid, the latter is
@@
Now, $p$ makes $\mathcal{W}$ into a $\overline{\mathcal{W}}$-module, and since $p^{L}$ is $\mathcal{V}$-linear and $\overline{\mathcal{W}}$ is rigid over $\mathcal{W}$, $p^{L}$ is $\overline{\mathcal{W}}$-linear by Section˜4.2. It follows then that $\hom^{\overline{\mathcal{W}}}_{\overline{\mathcal{W}}}(p^{L}(x),\mathbf{1}_{\overline{\mathcal{W}}})\simeq\hom^{\overline{\mathcal{W}}}_{\mathcal{W}}(x,p(\mathbf{1}_{\overline{\mathcal{W}}}))=\hom^{\overline{\mathcal{W}}}_{\mathcal{W}}(x,\mathbf{1}_{\mathcal{W}})=p^{R}\hom_{\mathcal{W}}^{\mathcal{W}}(x,\mathbf{1}_{\mathcal{W}})$ and so with $x=\mathbf{1}_{\mathcal{W}}$, we find $\hom_{\overline{\mathcal{W}}}^{\overline{\mathcal{W}}}(p^{L}(\mathbf{1}_{W}),\mathbf{1}_{\overline{\mathcal{W}}})=p^{R}(\mathbf{1}_{\mathcal{W}})$.
@@
The final thing we do before describing rigidifications in order to prove ˜4.67 is to prove that they exist.
@@
For this, we simply note that since the unit of $\mathcal{W}$ is $\kappa$-compact, any symmetric monoidal functor $\mathcal{W}_{0}\to\mathcal{W}$ has a $\kappa$-filtered colimit-preserving right adjoint by Section˜4.2. Thus we may replace the right hand side by $\operatorname{Fun}^{L,\otimes,\kappa}_{\mathcal{V}}(\mathcal{W}_{0},\mathcal{W})$.
Similarly, the unit of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ is atomic, and so by Section˜4.2, we may replace the left hand side by $\operatorname{Fun}^{iL,\otimes}_{\mathcal{V}}(\mathcal{W}_{0},\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa}))$.
@@
By Section˜4.3, it suffices to prove that $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ has a rigidification.
Consider the poset $P$ of full-sub-$\mathcal{V}$-algebras of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ that are rigid. By Section˜4.2, for $\mathcal{W}_{0}\in P$, the inclusion $\mathcal{W}_{0}\to\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ is an internal left adjoint and in particular preserves $\kappa$-compacts, and $\mathcal{W}_{0}$ is $\kappa$-compactly generated - thus $P$ is small : it is a subposet of the poset of subcategories of $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})^{\kappa}$.
@@
Let $\mathcal{W}_{0}:=\mathrm{colim}_{C\in P}C$, the colimit being taken in $\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}))$. By Section˜4.2, $\mathcal{W}_{0}$ is rigid, and it certainly has a map to $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$. Its image in the sense of Section˜4.3 is rigid (by that lemma) and contains $C$ for every $C\in P$, thus proving that it has a largest element, thereby concluding the proof. ∎
@@ 4.3.1 Over $\mathbf{Sp}$
In the same vein as for [RAM24, Section 4], while the existence of rigidification is good to know, for certain purposes it can be useful to have a more concrete description (for example, to prove ˜4.67). Our goal here is to draw inspiration from Section˜4.3.1 to construct the rigidification by hand in terms of trace-class maps.
@@
This follows from Section˜4.3.1: in a locally rigid category, every compact map is trace-class by Section˜4.1. ∎
@@
(Rigidity) We first prove that $\mathrm{Rig}_{\kappa}(\mathcal{V})$ is rigid. Our goal is to apply Section˜4.2. First, the unit is $y(\mathbf{1}_{\mathcal{V}})$ and hence compact.
@@
It follows from Section˜4.2 that $\mathrm{Rig}_{\kappa}(\mathcal{V})$ is rigid.
@@
are all equivalences. The first and last one are equivalences because $\mathcal{M}$ is rigid and the unit of both $\mathrm{Rig}_{\kappa}(\mathcal{V}),\mathrm{Ind}(\mathcal{V}^{\kappa})$ are compact, so that Section˜4.2 applies. For the middle one, we note that by Section˜4.3.1 and Section˜4.1, $\mathcal{M}$ is generated under colimits by $\mathbb{Q}_{\geq 0}$-indexed diagrams along trace-class maps, so it suffices to show that these land in $\mathrm{Rig}_{\kappa}(\mathcal{V})$ for any symmetric monoidal internal left adjoint $f:\mathcal{M}\to\mathrm{Ind}(\mathcal{V}^{\kappa})$.
@@
Now let $m\in\mathcal{M}$ be compactly exhaustible, so we write $m\simeq\mathrm{colim}_{\mathbb{N}}m_{n}$ with trace-class transition maps by Section˜4.1. Each of the $f(m_{n})\to f(m_{n+1})$ factors through a $y(z_{n}),z_{n}\in\mathcal{V}^{\kappa}$ by the above observation, where $f(m_{n})\to y(z_{n})$ is trace-class. Thus $f(m)\simeq\mathrm{colim}_{\mathbb{N}}y(z_{n})$ with each $z_{n}\to z_{n+1}$ trace-class.
@@
where the inclusion $1]\to[2]$ is the long edge, and where $\mathbb{Q}_{\geq 0}^{\delta}$ denotes the set $\mathbb{Q}_{\geq 0}$ viewed as a discrete category.
@@
$P^{\delta}\times Q_{0}$$P\times Q_{0}$$P^{\delta}\times Q_{1}$$P\times Q_{1}$}}
@@
As $P$ is linearly ordered, it is a filtered colimit of posets of the form $[n]$, and so we may assume without loss of generality that it is of this form. Since the pushouts $n]\coprod_{n=0}[1]\simeq~[n+1]$ are preserved by $(-)^{\simeq}$ (noting that $P^{\simeq}\simeq P^{\delta}$), we may also assume $P=[1]$.
@@
1. was proved in Section˜4.3.
2. If $\mathcal{V}$ is locally rigid, then compact maps are trace-class, and it follows that $\hat{y}:\mathcal{V}\to\mathrm{Ind}(\mathcal{V}^{\kappa})$ lands in $\mathrm{Rig}(\mathcal{V})$ by Section˜4.3.1 : indeed, compactly exhaustible objects generate $\mathcal{V}$ under colimits, and if $x=\mathrm{colim}_{\mathbb{Q}_{\geq 0}}x_{r}$ with compact transition maps, then $\hat{y}(x)=\mathrm{colim}_{\mathbb{Q}_{\geq 0}}y(x_{r})$, and each $y(x_{r})\to y(x_{s})$ is trace-class, as it was so in $\mathcal{V}$. In particular, $\hat{y}$ the desired fully faithful left adjoint.
Conversely, if there is such a left adjoint, then Section˜4.2 kicks in, showing that $\mathcal{V}$ is locally rigid. ∎
@@
Construction 4.86.
Let $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$. By Section˜4.2, $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})$ is rigid and hence the canonical symmetric monoidal colimit-preserving functor $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})\to\mathcal{V}$ factors through $\mathrm{Rig}(\mathcal{V})$. Beware that in general, the induced map $\mathrm{Ind}(\mathcal{V}^{\mathrm{dbl}})\to\mathrm{Rig}(\mathcal{V})$, while always fully faithful, is not in general an equivalence.
@@
Indeed consider, in $\prod_{i}\mathcal{V}_{i}$, a trace-class map $(v_{i})\to(w_{i})$. Each $v_{i}\to w_{i}$ is trace-class in $\mathcal{V}_{i}$, and hence compact ($\mathbf{1}_{\mathcal{V}_{i}}$ is compact) and hence factors through a compact ($\mathcal{V}_{i}$ is compactly generated) and hence through a dualizable ($\mathcal{V}_{i}$ is rigid). As dualizables in $\prod_{i}\mathcal{V}_{i}$ are detected pointwise, this implies that $(v_{i})\to(w_{i})$ factors through a dualizable object. Thus the result follows from the previous example.
More generally, as an application of this description of rigidification together with Section˜4.1, we obtain:
@@
Note that this forgetful functor exists by Section˜4.2. Before proving this, we will need a preliminary which is about trace-class maps in lax pullbacks. First, we analyze internal homs in lax pullbacks, by analyzing a simpler case.
@@
Since $y\to z$ is also trace-class in $C_{01}$, Section˜4.1 guarantees that there is a homotopy between the induced lifts for $gf$, and thus that $gf$ admits a trace-class witness in $C$, as was needed. ∎
@@
By Section˜4.3.1, the span $\overset{\rightarrow}{\mathrm{fib}}(p_{0})\to C_{01}\leftarrow\overset{\rightarrow}{\mathrm{fib}}(p_{1})$ fits the hypotheses of the previous corollary. It thus suffices to show that the projection of $f$ (resp. $g$) to $\overset{\rightarrow}{\mathrm{fib}}(p_{i})$ is trace-class.
@@
[…]
From there, using Section˜4.3.1 and the fact that $f,g$ commute with sequential colimits, we find that $\eta$ is an equivalence on $\omega_{1}$-compacts and thus on all objects because $f,g$ commute with all colimits. ∎
@@
Let $\mathcal{V}_{0}\xrightarrow{p_{0}}\mathcal{V}_{01}\xleftarrow{p_{1}}\mathcal{V}_{1}$ be span in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$, and let $\mathcal{V}$ be their pullback, $\overset{\rightarrow}{\mathcal{V}}$ be their lax pullback as in Section˜4.3.1. The canonical inclusion $\mathcal{V}\to\overset{\rightarrow}{\mathcal{V}}$ induces an equivalence on rigidification.
@@
We are finally equipped to give a proof of Section˜4.3.1:
Proof of Section˜4.3.1.
By definition of rigidity, this forgetful functor factors as a composite $\mathrm{CAlg}^{\mathrm{rig}}\to\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})\to(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}}$ where the second functor preserves limits for general reasons; thus we are really considering the first functor $\mathrm{CAlg}^{\mathrm{rig}}\to\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})$. Now, again by Section˜4.2 this is a full subcategory inclusion, so it suffices to prove that it is closed under limits.